17. What pressure will be exerted by 0.7 mol of a gas at 23 °C when contained in a vessel whose volume is 1.1 L?


1
Expert's answer
2021-04-25T03:49:20-0400

Ideal gas law PV = nRT

P= nRT/V= 0.7*0.082*296/1.1= 15.4atm

Hence the pressure of the gas is 15.4atm